# Deriving Causal Order from Single-Variable Interventions: Guarantees & Algorithm

### v1

@@ Abstract
Targeted and uniform interventions to a system are crucial for unveiling causal relationships. While several methods have been developed to leverage interventional data for causal structure learning, their practical application in real-world scenarios often remains challenging. Recent benchmark studies have highlighted these difficulties, even when large numbers of single-variable intervention samples are available. In this work, we demonstrate, both theoretically and empirically, that such datasets contain a wealth of causal information that can be effectively extracted under realistic assumptions about the data distribution. More specifically, we introduce the notion of interventional faithfulness, which relies on comparisons between the marginal distributions of each variable across observational and interventional settings, and we introduce a score on causal orders. Under this assumption, we are able to prove strong theoretical guarantees on the optimum of our score that also hold for large-scale settings. To empirically verify our theory, we introduce Intersort, an algorithm designed to infer the causal order from datasets containing large numbers of single-variable interventions by approximately optimizing our score. Intersort outperforms baselines (GIES, PC and EASE) on almost all simulated data settings replicating common benchmarks in the field. Our proposed novel approach to modeling interventional datasets thus offers a promising avenue for advancing causal inference, highlighting significant potential for further enhancements under realistic assumptions.
@@ 1 Introduction
Causal structure learning is pivotal for understanding complex systems, aiming to discover causal relationships from data. This field spans disciplines such as biology (Meinshausen et al., 2016; Yu et al., 2004), medicine (Farmer et al., 2018; Joffe et al., 2012; Feuerriegel et al., 2024), and social sciences Baum-Snow & Ferreira (2015); Imbens & Rubin (2015), where causal insights drive informed decision-making and deepen our understanding of underlying processes. Traditionally, causal discovery has relied heavily on observational data, primarily because expansive interventional experiments are often impractical or costly in many application domains. The inherent limitations of observational data necessitate assumptions about data distribution to ensure identifiability beyond the Markov equivalence class (Spirtes et al., 2000; Shimizu et al., 2006; Hoyer et al., 2008). However, the emergence of large-scale interventional data, especially in single-cell transcriptomics data (Replogle et al., 2022; Datlinger et al., 2017, 2021; Dixit et al., 2016), introduces both new opportunities and challenges.
Interventional data, characterized by targeted alterations to a system, offers a unique perspective for causal discovery. These controlled interventions reveal causal mechanisms often obscured in observational studies. As such, many causal discovery methods using interventional data have been proposed (Brouillard et al., 2020; Lorch et al., 2022; Lopez et al., 2022; Hauser & Bühlmann, 2012). However, their practicality remains challenging. For example, recent benchmarks in the gene network domain (Chevalley et al., 2023, 2022) demonstrate that simpler models based on straightforward statistical comparisons between observational and interventional distribution outperforms established causal discovery methods, particularly when many single-variable interventions are available. The motivation for this research thus stemmed from these recent advances in applying causal discovery techniques to gene network inference. Our methodology, including the use of the Wasserstein distance, aligns with evaluation metrics for predicted graphs as established in the benchmark by Chevalley et al. (2022). Additionally, a component of our algorithm builds upon and formalizes approaches from leading models in gene network inference, such as those highlighted in the CausalBench challenge (Kowiel et al., 2023; Nazaret & Hong, 2023; Deng & Guan, 2023; Chevalley et al., 2023). This work not only addresses a critical domain of application but also enhances the theoretical framework surrounding contemporary advancements in the field (Yu et al., 2004; Chai et al., 2014; Akers & Murali, 2021; Hu et al., 2020).
@@
Our empirical evaluations on diverse simulated datasets (linear, random Fourier features (Lorch et al., 2022), neural network (Nazaret et al., 2023; Brouillard et al., 2020) and single cell (Dibaeinia & Sinha, 2020) with various noise distributions), which are close replications of real-world systems and follow common practice in the field, confirm our theoretical results. The efficacy of our approach in accurately determining the causal order from interventional data signifies a notable advancement in causal inference. Intersort outperforms the three baselines, namely PC (Spirtes et al., 2000), GIES (Hauser & Bühlmann, 2012), and EASE (Gnecco et al., 2021) on almost all settings. It is also robust to data normalization (Reisach et al., 2021), which is a failure mode of many recently proposed continuous optimization causal discovery methods. Moreover, our empirical results show that Intersort makes more efficient use of interventional information compared to existing approaches which is a critical advantage especially in domains where interventions are costly and sometimes impossible to perform. This suggests that the task of identifying causal structures with many single-variable interventions may be more feasible than previously believed. Moreover, our method demonstrates similar performance and robustness across different types of simulated data, emphasizing its reliability and versatility.
@@ 2 Related work
Even though the causal order does not contain the full causal information as it does not uniquely identify the causal graph, it can subsequently be used with for example penalized regression techniques to recover the edges (Bühlmann et al., 2014; Shimizu et al., 2011). Also, the correct causal order is useful in itself as a fully connected graph can be constructed from it which describes the interventional distributions (Peters & Bühlmann, 2015; Bühlmann et al., 2014). More recently, following the discovery of Reisach et al. (2021) that the causal order can be recovered by sorting the variable by variance in simulated datasets, many algorithms to recover the causal order from observational data have been proposed, for example via score matching (Rolland et al., 2022; Montagna et al., 2023a, b). To our knowledge, we are the first to propose an algorithm to infer the causal order from interventional data.
@@ 3 Causal Order: Definitions
We consider a set of $d$ random variables $X=(X_{1},...,X_{d})$ indexed by $V=\{1,...d\}$, with associated joint distribution $P_{X}$. We denote the marginal distribution of each random variable as $P_{X_{i}}$, $i\in V$.
@@
To measure the discrepancy between a permutation $\pi$ and a graph $\mathcal{G}$, we use the top order divergence (Rolland et al., 2022). The top order divergence $D_{top}:(\{1,...d\},\{1,...d\}\times\{1,...d\})\times(\{1,...d\}\rightarrow\{1,...d\})\rightarrow\mathbb{N}_{0}$ is defined as:
@@
A structural causal model (SCM) $\mathcal{C}=(\textbf{S},P_{N})$ is determined by a collection of $d$ assignments $s\in\textbf{S}$
@@
where $\textbf{Pa}_{j}\in V\setminus j$ and $P_{N}$ is a joint distribution over the noise variables that is jointly independent (Peters et al., 2017).
Distribution with Intervention $P_{X}^{C,do(X_{k}:=\tilde{N}_{k}))}$
@@ 4 $\epsilon$-Interventional faithfulness
We here introduce a new concept that we name interventional faithfulness. Akin to the common assumption of faithfulness of a joint distribution with respect to the true graph, which assumes that all d-separations in the graph are reflected by conditional independences in the entailed distribution, $\epsilon$-interventional faithfulness assumes that all paths in the graphs are revealed by changes in distribution under intervention above a significance threshold $\epsilon$. This assumption is also similar to how Mooij et al. (2016) define causality as the existence of interventions inducing a change in distribution. We formalize our assumption as follows:
@@ 5 A score on causal orders
Given an observational distribution $P_{X}^{\mathcal{C},(\emptyset)}$ and a set of interventional distributions $\mathcal{P}_{int}=\{P_{X}^{\mathcal{C},do(X_{k}:=\tilde{N}_{k})},k\in\mathcal{I}\}$, $\mathcal{I}\subseteq V$, we define the following score for a permutation $\pi$, for some statistical distance $D:\mathcal{P}(M)\times\mathcal{P}(M)\rightarrow[0,\infty)$, $\epsilon>0$:
$\mathcal{S}(\pi,\epsilon,D,\mathcal{I},P_{X}^{\mathcal{C},(\emptyset)},\mathcal{P}_{int})=\sum_{\pi(i)<\pi(j),i\in\mathcal{I},j\in V}\left(D\left(P_{X_{j}}^{\mathcal{C},(\emptyset)},P_{X_{j}}^{\mathcal{C},do(X_{i}:=\tilde{N}_{i})}\right)-\epsilon\right)$ (1)
Usually, we can set the $\epsilon$ of eq. 1 to the same value of $\epsilon$-interventional faithfulness as the input dataset, unless it is equal to zero. However, the remark in remark Remark ensures the existence of a suitable $\epsilon$ in that case, as it can be chosen in a data-driven way by taking a non-zero value for $\epsilon$ smaller than the smallest non-negative distance.
@@
The above theorem shows that in the case where we observe interventions on all the variables, we are guaranteed to find a valid topological order by maximizing our score when the data is $\epsilon$-interventionally faithful. We now turn to the case where variables have a probability of being intervened smaller than one. In that case, the error will not be zero in expectation, and we thus aim to upper-bound the expected error as measured by $D_{top}$. We first prove an important property of the optimum to our proposed score. Lemma 4 shows that for a given edge, a single intervention among a large set of candidate variables is sufficient to correctly order the two variables of an edge.
@@ 7 Intersort
Given that our score for causal order eq. 1 is optimized over a set of permutations, finding the optimum by exhaustively trying all possible solutions becomes intractable as the number of variables grows larger. We thus develop an approximation algorithm to tractably find a solution that optimizes for our proposed score in eq. 1. This algorithm consists of two steps. The first step uses a different score to find an initial solution. In the second step, this solution is refined by performing a local search using our original score eq. 1.
@@ 7.2 Intersort step 2: A local search for $\pi_{opt}$
Algorithm 1 Complete algorithm
1:procedure Intersort( $\mathcal{S}$: score function eq. 1, $\epsilon$,
$D$: statistical distance function,
$\mathcal{I}$: index set of intervened variables,
$P_{X}^{\mathcal{C},(\emptyset)}$: observational distribution,
$\mathcal{P}_{int}$: interventional distributions )
2:    Initialize $\mathbf{D}$ as a zero matrix of size $d\times d$
3:    for each $(i,j)\in[1..d]$ do
4:         if $i\in\mathcal{I}$ then
6:         end if
7:    end for
8:    $\pi_{init}\leftarrow\textsc{sortranking}(\mathbf{D},\epsilon)$
10:    return $\pi_{opt}$
11:end procedure
Starting from the initial solution found by maximizing eq. 2, we then perform a greedy local search to optimize our original objective eq. 1. At each step, given a current solution, we search in a close neighbourhood set for a candidate permutation $\pi$ with a higher score than our current solution. To derive the neighborhood set, we introduce an operator $\Sigma:\{\{1,...d\}\rightarrow\{1,...d\}\}\rightarrow\{\{1,...d\}\rightarrow\{1,...d\}\}$, defined such that for a given set of permutations $\Pi$, $\Sigma(\Pi)$ returns a new set of permutations $\Pi^{\prime}$. Each permutation $\pi\in\Pi$ is modified by applying $\sigma:\{1,...d\}\rightarrow\{1,...d\}\rightarrow\{\{1,...d\}\rightarrow\{1,...d\}\}$, which for each variable $i\in V$, returns all the permutations such that $\pi(i)$ takes any of the other $d-1$ possible values. $\sigma$ takes one permutation as input and return a set of permutations which is $\mathcal{O}(d^{2})$. The total number of applications of $\Sigma$, denoted as $k$, directly defines the neighborhood size which is $\mathcal{O}(d^{2k})$. This definition allows us to explore varying depths of the search space, which is $\mathcal{O}(d!)$, by controlling $k$, thereby balancing between the breadth of the search and the computational resources required. A smaller $k$ value implies a tighter search around the initial solution and encode our confidence in the initial solution, optimizing runtime at the potential cost of missing broader, potentially better solutions. In our experiments, we use $k=1$. We then stop once an iteration of the algorithm brings no improvement to the score. This procedure, that we name localsearch, is described in algorithm 3 in the appendix. Our complete algorithm for finding the causal order is called Intersort (see algorithm 1).
@@
We note that more refined algorithms could be developed to solve this optimization over permutations. For example, for large $d$, even the local search may become computationally expensive. However, for the purpose of this study, we find the proposed algorithm to have satisfying properties, as it scales to settings with $30$ variables and finds a solution close to the optimum. As such, we leave further algorithmic development for large scale settings as potential future work.

### v3

@@ Abstract
Targeted and uniform interventions to a system are crucial for unveiling causal relationships. While several methods have been developed to leverage interventional data for causal structure learning, their practical application in real-world scenarios often remains challenging. Recent benchmark studies have highlighted these difficulties, even when large numbers of single-variable intervention samples are available. In this work, we demonstrate, both theoretically and empirically, that such datasets contain a wealth of causal information that can be effectively extracted under realistic assumptions about the data distribution. More specifically, we introduce a novel variant of interventional faithfulness, which relies on comparisons between the marginal distributions of each variable across observational and interventional settings, and we introduce a score on causal orders. Under this assumption, we are able to prove strong theoretical guarantees on the optimum of our score that also hold for large-scale settings. To empirically verify our theory, we introduce Intersort, an algorithm designed to infer the causal order from datasets containing large numbers of single-variable interventions by approximately optimizing our score. Intersort outperforms baselines (GIES, DCDI, PC and EASE) on almost all simulated data settings replicating common benchmarks in the field. Our proposed novel approach to modeling interventional datasets thus offers a promising avenue for advancing causal inference, highlighting significant potential for further enhancements under realistic assumptions.
00footnotetext: Correspondence to m.chevalley97@gmail.com, arash@distantvantagepoint.com
@@ 1 Introduction
Causal structure learning is pivotal for understanding complex systems, aiming to discover causal relationships from data. This field spans disciplines such as biology (Meinshausen et al., 2016; Yu et al., 2004), medicine (Farmer et al., 2018; Joffe et al., 2012; Feuerriegel et al., 2024), and social sciences Baum-Snow & Ferreira (2015); Imbens & Rubin (2015), where causal insights drive informed decision-making and deepen our understanding of underlying processes. Traditionally, causal discovery has relied heavily on observational data, primarily because expansive interventional experiments are often impractical or costly in many application domains. The inherent limitations of observational data necessitate assumptions about data distribution to ensure identifiability beyond the Markov equivalence class (Spirtes et al., 2000; Shimizu et al., 2006; Hoyer et al., 2008). However, the emergence of large-scale interventional data, especially in single-cell transcriptomics data (Replogle et al., 2022; Datlinger et al., 2017; 2021; Dixit et al., 2016), introduces both new opportunities and challenges.
Interventional data, characterized by targeted alterations to a system, offers a unique perspective for causal discovery. These controlled interventions reveal causal mechanisms often obscured in observational studies. As such, many causal discovery methods using interventional data have been proposed (Brouillard et al., 2020; Lorch et al., 2022; Lopez et al., 2022; Hauser & Bühlmann, 2012). However, their practicality remains challenging. For example, recent benchmarks in the gene network domain (Chevalley et al., 2023; 2022) demonstrate that simpler models based on straightforward statistical comparisons between observational and interventional distribution outperforms established causal discovery methods, particularly when many single-variable interventions are available. The motivation for this research thus stemmed from these recent advances in applying causal discovery techniques to gene network inference. Our methodology, including the use of the Wasserstein distance, aligns with evaluation metrics for predicted graphs as established in the benchmark by Chevalley et al. (2022). Additionally, a component of our algorithm builds upon and formalizes approaches from leading models in gene network inference, such as those highlighted in the CausalBench challenge (Kowiel et al., 2023; Nazaret & Hong, 2023; Deng & Guan, 2023; Chevalley et al., 2023). This work not only addresses a critical domain of application but also enhances the theoretical framework surrounding contemporary advancements in the field (Yu et al., 2004; Chai et al., 2014; Akers & Murali, 2021; Hu et al., 2020).
@@
Our empirical evaluations on diverse simulated datasets (linear, random Fourier features (Lorch et al., 2022), neural network (Nazaret et al., 2023; Brouillard et al., 2020) and single cell (Dibaeinia & Sinha, 2020) with various noise distributions), which are close replications of real-world systems and follow common practice in the field, confirm our theoretical results. The efficacy of our approach in accurately determining the causal order from interventional data signifies a notable advancement in causal inference. Intersort outperforms the four baselines, namely PC (Spirtes et al., 2000), GIES (Hauser & Bühlmann, 2012), DCDI (Brouillard et al., 2020) and EASE (Gnecco et al., 2021) on almost all settings. It is also robust to data normalization (Reisach et al., 2021), which is a failure mode of many recently proposed continuous optimization causal discovery methods. Moreover, our empirical results show that Intersort makes more efficient use of interventional information compared to existing approaches which is a critical advantage especially in domains where interventions are costly and sometimes impossible to perform. This suggests that the task of identifying causal structures with many single-variable interventions may be more feasible than previously believed. Moreover, our method demonstrates similar performance and robustness across different types of simulated data, emphasizing its reliability and versatility.
@@ 2 Related work
Even though the causal order does not contain the full causal information as it does not uniquely identify the causal graph, it can subsequently be used with for example penalized regression techniques to recover the edges (Bühlmann et al., 2014; Shimizu et al., 2011). Also, the correct causal order is useful in itself as a fully connected graph can be constructed from it which describes the interventional distributions (Peters & Bühlmann, 2015; Bühlmann et al., 2014). More recently, following the discovery of Reisach et al. (2021) that identified an artifact in simulated datasets where sorting variables by variance appears to recover the causal order under certain conditions, many algorithms to recover the causal order from observational data have been proposed, for example via score matching (Rolland et al., 2022; Montagna et al., 2023a; b). To our knowledge, another work aiming to infer the causal order from interventions comes from Tian & Pearl (2001), which builds on seminal works such as Cooper & Yoo (1999); Spirtes et al. (2000). Tian & Pearl (2001) focus on causal discovery from passive interventions, where shifts in distributions arise naturally. Our work brings two major contributions: first, we propose a score-based approach, in contrast to the rule-based algorithm of Tian & Pearl (2001), which opens the door for a large class of optimization tools and offers better scalability, and second, we provide extended theoretical results in terms of upper-bounding the expected error of our algorithm, especially in cases where only a subset of the variables is intervened on whereas Tian & Pearl (2001) argues mainly about the recovery of the true causal order when ”all” variables are intervened. Adjacently, methods like DirectLiNGAM (Shimizu et al., 2011) infer causal orderings from observational data under non-Gaussianity assumptions. As such, they score orderings based on fitted parameters, making it sensitive to distributional assumptions. By leveraging interventional data, our approach relaxes distributional assumptions and provides a more robust framework for causal ordering. Eberhardt et al. (2005; 2006) explored the theoretical bounds on the number of experiments required for full causal structure identification. Our work complements these findings by focusing on deriving causal orderings from single-variable interventions under light assumptions, which can be more practical in large-scale applications.
@@ 3 Causal Order: Definitions
We consider a set of $d$ random variables $X=(X_{1},...,X_{d})$ indexed by $V=\{1,...,d\}$, with associated joint distribution $P_{X}$. We denote the marginal distribution of each random variable as $P_{X_{i}}$, $i\in V$.
@@
To measure the discrepancy between a permutation $\pi$ and a graph $\mathcal{G}$, we use the top order divergence (Rolland et al., 2022). The top order divergence $D_{top}:(\{1,...,d\},\{1,...,d\}\times\{1,...,d\})\times(\{1,...,d\}\rightarrow\{1,...,d\})\rightarrow\mathbb{N}_{0}$ is defined as:
@@
An acyclic structural causal model (SCM) $\mathcal{C}=(\textbf{S},P_{N})$ is determined by a collection of $d$ assignments $s_{j}\in\textbf{S}$, $\mid\textbf{S}\mid=d$, such that for each assignment $s_{j}$ defines $X_{j}$
@@
where $\textbf{Pa}_{j}\in V\setminus j$ is the set of variables on which $X_{j}$ directly depends on and $P_{N}$ is a joint distribution over the noise variables $N=\left(N_{1},...,N_{d}\right)$ that is jointly independent (Peters et al., 2017). The associated graph $\mathcal{G}$ that connects variables to their parents is assumed to be acyclic.
Distribution with Intervention $P_{X}^{(C,do(X_{k}:=\tilde{N}_{k}))}$
@@ 4 $\epsilon$-Interventional faithfulness
We here introduce the main assumption of our work that we name interventional faithfulness. Akin to the common assumption of faithfulness of a joint distribution with respect to the true graph, which assumes that all d-separations in the graph are reflected by conditional independences in the entailed distribution, $\epsilon$-interventional faithfulness assumes that all paths in the graphs are revealed by changes in distribution under intervention above a significance threshold $\epsilon$. This assumption is also similar to how Mooij et al. (2016) define causality as the existence of interventions inducing a change in distribution. The main difference with comparable assumptions made in the literature is that this definition depends on a statistical distance and significance threshold, instead of an oracle of the statistical dependence. We formalize our assumption as follows:
@@
The concept of interventional faithfulness, where interventions on a node cause significant deviations in the distribution of downstream nodes, is discussed in various contexts. It aligns closely with the notion of ”c-faithfulness” in causal graphs, where interventional distributions respect the causal structure and result in significant changes in downstream nodes (Shanmugam et al., 2015; Squires et al., 2020), and with the notion of ”influentiality” (Tian & Pearl (2001), Definition 2). Assumption 4.4 in Yang et al. (2018) explicitly states that interventions on upstream nodes should affect downstream nodes. The causal Markov condition and faithfulness for both observational and interventional distributions also support this idea (Addanki & Kasiviswanathan, 2021). The use of interventional data to improve the identifiability of causal models is another related concept, emphasizing how interventions can reveal causal relationships not identifiable from observational data alone (Hauser & Bühlmann, 2014).
@@
5 Latent confounders
Until now, we have assumed that all causal variables are observed, an assumption called causal sufficiency. We here discuss how the notion of interventional faithfulness and the results of this paper are applicable to the case where only a subset of the causal variables are observed. This is a common situation in practice either because we want to focus on a subset of variables which are of our interest, or because we are certain about the existence of unobserved confounders. To that end, we take the common formalization of marginalized SCMs, that marginalizes the latent confounders from the joint distribution and gives rise to a new SCM where the effects of marginalized variables are now reflected in functional dependencies among observed variables. Therefore, some variables become confounded as they depend on the same noise variables. An important property of acyclic SCMs is that they are closed under marginalizations and that their marginalization respects the latent projection (Bongers et al., 2021). The latent projection describes how to construct a new graph among the remaining variables, representing the appearance of confounders, or dependent noise variables, with bidirected (hyper) edges (Verma, 1991; Evans, 2016; 2018b). Those two representations are referred to as Acyclic Directed Mixed Graphs (ADMGs) (Verma, 1991; Andrews et al., 2022; Verma & Pearl, 2022; Richardson, 2003) and mDAGs (Evans, 2016; 2018a; Bongers et al., 2021). For our purpose, both representations are equivalent, and we thus focus on ADMGs. For the DAG $\mathcal{G}$ of the original SCM, we write the projection over a subset of variables $V^{\prime}\subset V$ as $\mathcal{G}^{\prime}=(V^{\prime},E^{\prime},B)$, where $B$ is a set of bidirected edges, and the marginalized SCM as $\mathcal{C}^{\prime}$.
Proposition 1.
If the tuple $(\tilde{N},\mathcal{C})$ is $\epsilon$-interventionally faithful to the graph $\mathcal{G}$ associated to $\mathcal{C}$, then for any marginalization $V^{\prime}\subset V$, the marginalized SCM $\mathcal{C}^{\prime}$ and the subset of interventional distributions $\tilde{N}^{\prime}=\bigcup_{k\in\mathcal{I}\bigcap V^{\prime}}\{\tilde{N}_{k}\}$ is $\epsilon$-interventionally faithful to $(V^{\prime},E^{\prime})$ from the projected ADMG $\mathcal{G}^{\prime}=(V^{\prime},E^{\prime},B)$.
This result derives directly from the fact that ancestral relationships are preserved in the projected graph, and that the (interventional) distributions induced by the marginalized SCM correspond to the marginals of the distributions induced by the original SCM (Bongers et al., 2021). As such, our algorithm is applicable to settings with confounders or on subsets of variables, and the theoretical guarantees hold for the projected DAG $(V^{\prime},E^{\prime})$. However, we note that we here implicitly assume that there is no selection bias in the sampling process (e.g., preferential inclusion of certain data points or experiments), and that latent confounders remain stable across conditions.
@@ 6 A score on causal orders
Given an observational distribution $P_{X}^{\mathcal{C},(\emptyset)}$ and a set of interventional distributions $\mathcal{P}_{int}=\{P_{X}^{\mathcal{C},do(X_{k}:=\tilde{N}_{k})},k\in\mathcal{I}\}$, $\mathcal{I}\subseteq V$, we define the following score for a permutation $\pi$, for some statistical distance $D:\mathcal{P}(M)\times\mathcal{P}(M)\rightarrow[0,\infty)$, $\epsilon>0$, $c>\epsilon$:
$\begin{split}S(\pi,\epsilon,D,\mathcal{I},P_{X}^{\mathcal{C},(\emptyset)},\mathcal{P}_{int},c)=&\sum_{\pi(i)<\pi(j),i\in\mathcal{I},j\in V}\left(D\left(P_{X_{j}}^{\mathcal{C},(\emptyset)},P_{X_{j}}^{\mathcal{C},do(X_{i}:=\tilde{N}_{i})}\right)-\epsilon\right)\\ &+c\cdot d\cdot\bm{1}_{D\left(P_{X_{j}}^{\mathcal{C},(\emptyset)},P_{X_{j}}^{\mathcal{C},do(X_{i}:=\tilde{N}_{i})}\right)>\epsilon}\end{split}$ (1)
Intuitively, the sum quantifies how the causal order corresponds to strong causal effects. The rescaling from the second term by a factor of $d$ ensures that effects higher than $\epsilon$ will force the optimal solution to have $\pi(i)<\pi(j)$ by inflating their weight compared to effects smaller than $\epsilon$. Usually, we can set the $\epsilon$ of eq. 1 to the same value of $\epsilon$-interventional faithfulness as the input dataset, unless it is equal to zero. However, the remark in remark Remark ensures the existence of a suitable $\epsilon$ in that case, as it can be chosen in a data-driven way by taking a non-zero value for $\epsilon$ smaller than the smallest non-negative distance.
@@
The above theorem shows that in the case where we observe interventions on all the variables, we are guaranteed to find a valid topological order by maximizing our score when the data is $\epsilon$-interventionally faithful. Similar results were proven for this case where all the variables are intervened (Hauser & Bühlmann, 2012; Tian & Pearl, 2001). We now turn to the case where variables have a probability of being intervened smaller than one. In that case, the error will not be zero in expectation, and we thus aim to upper-bound the expected error as measured by $D_{top}$. We first prove an important property of the optimum to our proposed score. Lemma 4 shows that for a given edge, a single intervention among a large set of candidate variables is sufficient to correctly order the two variables of an edge.
@@ 8 Intersort
Given that our score for causal order eq. 1 is optimized over a set of permutations, finding the optimum by exhaustively trying all possible solutions becomes intractable as the number of variables grows larger. We thus develop an algorithm to tractably find a solution that optimizes for our proposed score in eq. 1. This algorithm consists of two steps. The first step uses a different score to find an initial solution. In the second step, this solution is refined by performing a local search using our original score eq. 1.
@@ 8.2 Intersort step 2: A local search for $\pi_{opt}$
Starting from the initial solution found by maximizing eq. 2, we then perform a greedy local search to optimize our original objective eq. 1. At each step, given a current solution, we search in a close neighbourhood set for a candidate permutation $\pi$ with a higher score than our current solution. To derive the neighborhood set, we introduce an operator $\Sigma:\{\{1,...d\}\rightarrow\{1,...d\}\}\rightarrow\{\{1,...d\}\rightarrow\{1,...d\}\}$, defined such that for a given set of permutations $\Pi$, $\Sigma(\Pi)$ returns a new set of permutations $\Pi^{\prime}$. Each permutation $\pi\in\Pi$ is modified by applying $\sigma:\{1,...d\}\rightarrow\{1,...d\}\rightarrow\{\{1,...d\}\rightarrow\{1,...d\}\}$, which for each variable $i\in V$, returns all the permutations such that $\pi(i)$ takes any of the other $d-1$ possible values. $\sigma$ takes one permutation as input and return a set of permutations which is $\mathcal{O}(d^{2})$. The total number of applications of $\Sigma$, denoted as $k$, directly defines the neighborhood size which is $\mathcal{O}(d^{2k})$. This definition allows us to explore varying depths of the search space, which is $\mathcal{O}(d!)$, by controlling $k$, thereby balancing between the breadth of the search and the computational resources required. A smaller $k$ value implies a tighter search around the initial solution and encode our confidence in the initial solution, optimizing runtime at the potential cost of missing broader, potentially better solutions. In our experiments, we use $k=1$. We then stop once an iteration of the algorithm brings no improvement to the score. This procedure, that we name localsearch, is described in algorithm 3 in the appendix. Our complete algorithm for finding the causal order is called Intersort (see algorithm 1 in the appendix).
@@
We note that more refined algorithms could be developed to solve this optimization over permutations. For example, for large $d$, the local search may become computationally expensive. However, for the purpose of this study, we find the proposed algorithm to have satisfying properties, as it scales to settings with $30$ variables and finds a solution close to the optimum. We also report results for a scale-free network modelled with the Barabasi-Albert distribution Albert & Barabási (2002) in the appendix in fig. 3, which supports the generality of our approach. We also report the performance of SORTRANKING in large scale settings in figs. 5 and 4. The results demonstrate that the our score can be optimized at scale, even though we can observe room for improvement, as the error is above the upper-bounds for many settings. As such, we leave further algorithmic development for large scale settings as potential future work.